Question 21

Consider the information provided in question 20. What is the probability that the sample mean is at least 67.5?

(1) 2.5
(2) 0.9938
(3) 0.0062
(4) -2.5
(5) None of the above
1
Expert's answer
2020-03-03T17:44:04-0500

2)0.9938

P(X"\geq" 67.5)

Mean=70, sd=6, sample size =36

"Z=\frac {(X-\mu) *\sqrt n} {SD}"

"Z=\frac{(67.5-70)*\sqrt{36}}{6}"

=-2.5

"\Phi(2.5)"

The value is 0.9938 from the z-tables. The answer is (2) 0.9938.